OpenStudy (anonymous):

If f(x) = x2, find f(-1).

OpenStudy (anonymous):

\[f(x) = x^2 \rightarrow f(-1) =(-1)^2 \rightarrow f(-1) =(-1)^2\] @kcurri

OpenStudy (anonymous):

just substitude

OpenStudy (anonymous):

f(-1)=(-1)*2

OpenStudy (anonymous):

substitude what?

OpenStudy (anonymous):

so its -2

OpenStudy (anonymous):

wait the (2) is an exponent??

OpenStudy (anonymous):

if yes it will be 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np no problem hv a nice day:)

OpenStudy (anonymous):

You to! :)